# Real-Time Neural Video Compression with Unified Intra and Inter Coding

### v4

@@ 3.5 Training with hybrid references
The core challenge in training UI2C is enabling the model to dynamically balance intra- and inter-coding based on the current reference error level. For the reference of initial frames, we consider three candidates: a pure blank signal (e.g., an all-zero image), the ground-truth (GT) of the previous frame, and a noise-corrupted version of this GT. During training, we randomly sample one of these three as the initial frame’s reference. This strategy forces the model to implicitly learn to assess reference error levels—allowing it to adaptively enhance intra-coding for error correction when processing sequences longer than training data, without manual reference discarding. Eliminating the need for info-discarding refresh mechanisms also reduces peak bitrate, thereby mitigating the risk of network congestion.
@@
    Dec. step: The number of autoregressive steps during the decoding of latent variables per frame (excluding optical flow and hyperprior) in the entropy model.
@@
Figure 6: Bitrate and quality variation across frames. The tested videos are Cactus, BQTerrace, and Bosphorus, from left to right. We compare two settings: divided (intra and inter frames use two models, refresh period is 64) and unified (all frames use the same model, no refresh). Note that we test single-frame compression models in this figure, rather than the proposed two-frame compression scheme. It is evident that the unified model delivers more stable bitrate and quality per frame and avoids the need for refresh, even in the setting of single-frame compression. Yet, our two-frame compression scheme performs even better.
@@ 4.1 Setting
Training Details. To achieve multi-rate, we randomly selected different quantization parameters (QPs) in the range [0, 63] for each training iteration. In a group of 8 images, the QP bias selection was [0, 8, 0, 4, 0, 4, 0, 4] for hierarchical quality. We followed [21] to assign different hierarchical weights to the loss function of each frame to support the hierarchical quality structure. For the loss function, we used a scaled YUV mean squared error (MSE) loss. The training was conducted on 8 NVIDIA RTX 4090 GPUs.
@@ 4.2 Comparison results
Figure 4 illustrates the rate-distortion curves on the HEVC-E and HEVC-B test sets. Our model consistently outperforms VTM across almost all quality levels. In particular, it shows strong performance at low bitrates. At high bitrates, thanks to reduced error accumulation in long sequences, our model even surpasses the highly complex DCVC-FM on HEVC-E. However, our model does not perform as well on shorter sequences, especially on MCL-JCV, which contains sequences with a maximum length of only 150 frames. Additionally, further exploration of training strategies is still needed—for instance, our reproduced version of DCVC-RT still trails the official model by about 20.7%. Nevertheless, under almost similar training settings, our model already outperforms the released version of DCVC-RT.
Moreover, our model demonstrates a more stable rate and quality control, and recovers more quickly to higher quality after scene changes, as shown in Figure 1. We tested on the kimono sequence, which contains scene cuts, and observed that our model regains quality faster compared to DCVC-RT, indicating its stronger capability in handling scenes with abrupt content changes. Moreover, our approach maintains consistently high quality without requiring refresh operations and remains unaffected by error propagation. Moreover, while sustaining higher quality, our method achieves better bitrate savings, and the peak bitrate has also been reduced. Additional sequence results are provided in the Appendix.
In Figure 5, we compare the intra-coding performance of our model with that of DCVC-RT. It can be observed that the intra-frame compression ability of our model is significantly better than that of DCVC-RT’s P-frame model, and is only slightly worse than DCVC-RT’s high-complexity I-frame model. This experiment suggests that intra- and inter-frame coding can be effectively unified within a single model. Our model opens up possibilities for further exploration of unified NVC architectures.
@@ 4.3 Complexity analysis
Table 2 provides a complexity comparison among different models. Compared to DCVC-DC and DCVC-FM, our model achieves a better compression ratio while maintaining lower computational complexity. Relative to DCVC-RT, our approach exhibits slightly higher encoding and decoding complexity. However, since our method processes two frames jointly during encoding and decoding, the average latent size per frame and the number of decoding steps are reduced by half. As a result, the overall frame rates of our method and DCVC-RT are comparable. It should be noted that our method introduces higher latency due to the one-frame delay in the encoding process. This trade-off is acceptable in scenarios where low latency is not strictly required.
@@ 4.4 Ablation studies
Unified Intra- and Inter-Frame Coding. To evaluate the efficacy of unifying intra- and inter-frame coding within a single model, we perform a comparative experiment. Specifically, for the baseline without this technique, we employ the I-frame model from RT to handle intra-frame coding. Experimental results demonstrate that when an I-frame model is available, the refresh mechanism proposed by DCVC-FM leads to significant performance gains. However, in the absence of the refresh mechanism, as shown in Table 3, severe error accumulation occurs, resulting in substantially degraded performance.
Meanwhile, we explored a unified model for both intra-frame and inter-frame processing within a single frame. Experiments show that after enhancing the intra-frame capability, the model’s performance improved by 64.9% under the non-refresh IP-1 condition. However, under the refresh condition, the performance was inferior to the non-refresh scenario, as the model tends to lose information and was not optimized using three-channel images of I-frames during training, while the unified model already handles error propagation effectively.
@@
Simultaneous Two-Frame Compression. To assess the effectiveness of the two-frame compression technique, we compare it against a model configuration without any frame delay. As presented in Table 3, the introduction of the two-frame compression technique yields a remarkable performance improvement, even in scenarios without any refresh operation.

### v5

@@ 3.5 Training with hybrid references
The core challenge in training UI2C is enabling the model to dynamically balance intra- and inter-coding based on the current reference error level. For the reference of initial frames, we consider three candidates: a pure blank signal (e.g., an all-zero image), the ground-truth (GT) of the previous frame, and a noise-corrupted version of this GT. For the noisy-reference features, we infer them from previously reserved frames and use the resulting features as training references. During training, we randomly sample one of these three as the initial frame’s reference. This strategy forces the model to learn to implicitly assess reference error levels, allowing it to adaptively enhance intra-coding for error correction when processing sequences longer than the training data, without requiring manual reference discarding. Eliminating the need for info-discarding refresh mechanisms also reduces peak bitrate, thereby mitigating the risk of network congestion.
@@
    Dec. step: The number of average autoregressive steps during the decoding of latent variables per frame (excluding optical flow and hyperprior) in the entropy model.
@@
Figure 6: Bitrate and quality variation across frames. The tested videos are Cactus, BQTerrace, and Bosphorus, from left to right. We compare two settings: divided (intra and inter frames use two models, refresh period is 64) and unified (all frames use the same model, no refresh). Note that we test single-frame compression models in this figure, rather than the proposed two-frame compression scheme. It is evident that the unified model delivers more stable bitrate and quality per frame and avoids the need for refresh, even in the setting of single-frame compression. Yet, our two-frame compression scheme performs even better. Detailed results for more sequences are provided in the supplementary material.
@@ 4.1 Setting
Training Details. To achieve multi-rate, we randomly selected different quantization parameters (QPs) in the range [0, 63] for each training iteration. In a group of 8 images, the QP bias selection was [0, 8, 0, 4, 0, 4, 0, 4] for hierarchical quality. We followed [21] to assign different hierarchical weights to the loss function of each frame to support the hierarchical quality structure. For the loss function, we used a scaled YUV mean squared error loss. The training was conducted on 8 NVIDIA RTX 4090 GPUs.
@@ 4.2 Comparison results
Figure 4 illustrates the rate-distortion curves on the HEVC-E and HEVC-B test sets. Our model consistently outperforms VTM across almost all quality levels. In particular, it shows strong performance at low bitrates. At high bitrates, thanks to reduced error accumulation in long sequences, our model even surpasses the highly complex DCVC-FM on HEVC-E. However, our model does not perform as well on shorter sequences, especially on MCL-JCV, which contains sequences with a maximum length of only 150 frames. Additionally, further exploration of training strategies is still needed—for instance, our reproduced version of DCVC-RT still trails the official model. Nevertheless, under almost similar training settings, our model already outperforms the released version of DCVC-RT.
Moreover, our model exhibits more stable bitrate and quality fluctuations, and recovers more rapidly to higher quality after scene changes, as shown in Figure 1. Evaluated on the kimono sequence, which contains explicit scene cuts, our model restores quality noticeably faster than DCVC-RT, demonstrating a stronger ability to cope with abrupt content transitions. Moreover, our approach maintains consistently high quality without requiring refresh operations and remains unaffected by error propagation. Moreover, while sustaining higher quality, our method achieves better bitrate savings, and the peak bitrate has also been reduced. Detailed results for more sequences are provided in the supplementary material.
In Figure 5, we compare the intra-coding performance of our model with that of DCVC-RT. It can be observed that the intra-frame compression ability of our model is significantly better than that of DCVC-RT’s P-frame model, and is only slightly worse than DCVC-RT’s high-complexity I-frame model. This experiment suggests that our UI2C’s intrinsic intra-coding capacity is markedly superior to that of DCVC-RT’s P-frame model, allowing it to handle previously unseen scenes more effectively.
@@ 4.3 Complexity analysis
Table 2 provides a complexity comparison among different models. Compared to DCVC-DC and DCVC-FM, our model achieves a better compression ratio while maintaining lower computational complexity. Relative to DCVC-RT, our approach exhibits slightly higher encoding and decoding complexity. However, since our method processes two frames jointly during encoding and decoding, the average latent size per frame and the number of decoding steps are reduced by half. The overall frame rates of our method and DCVC-RT are comparable. It should be noted that our method introduces higher latency due to the one-frame delay in the encoding process. This trade-off is acceptable in scenarios where low latency is not strictly required.
@@ 4.4 Ablation studies
Unified Intra- and Inter-Frame Coding. To evaluate the efficacy of unifying intra- and inter-frame coding within a single model, we perform a comparative experiment. Specifically, for the baseline without this technique, we employ the I-frame model from RT to handle intra-frame coding. Table  3 (Rows 1 vs. 5) demonstrates that when an I-frame model is available, the refresh mechanism proposed by DCVC-FM leads to significant performance gains. However, in the absence of the refresh mechanism, severe error accumulation occurs, resulting in substantially degraded performance.
Meanwhile, we explored a unified model for both intra-frame and inter-frame processing within a single frame. Experiments show that after enhancing the intra-frame capability, the model’s performance improved by 64.9% under the non-refresh IP-1 condition (Table  3, Rows 5 vs. 6). However, under the refresh condition, the performance was inferior to the non-refresh scenario (Table  3, Rows 2 vs. 6), as the model tends to lose rich temporal feature information and was not optimized using three-channel images of I-frames during training, while the unified model already handles error propagation effectively.
@@
Simultaneous Two-Frame Compression. To assess the effectiveness of the two-frame compression technique, we compare it against the model configuration without any frame delay. As presented in Table 3, the introduction of the two-frame compression technique yields a remarkable performance improvement, even in scenarios without any refresh operation.